# Doubly Robust Alignment for Large Language Models

### v1

@@ 1 Introduction
Recent advances in large language models (LLMs) have revolutionized various natural language processing tasks, ranging from text generation to human-AI conversation and more complex reasoning tasks [Brown et al., , 2020, Touvron et al., , 2023, Guo et al., , 2025]. These models are typically trained in two stages. In the pre-training stage, LLMs learn general linguistic patterns and commonsense knowledge from vast, unlabeled text data through autoregressive next-token prediction. However, pretrained models face a critical objective mismatch: while they are optimized for token prediction, real-world deployment requires alignment with complex human values such as helpfulness, honesty and harmlessness [Askell et al., , 2021]. This mismatch calls for an additional post-training stage, aiming at better aligning these pre-trained models with human preference.
The paper studies reinforcement learning from human feedback (RLHF), a post-training paradigm that adapts pre-trained models through reinforcement learning [RL, Sutton et al., , 2018]. The RLHF literature has rapidly expanded in recent years, where existing algorithms can be broadly categorized as reward-based or preference-based (Section 2 for a review). While demonstrating remarkable success in domains including robotics control, video games, and LLMs fine-tuning [see e.g., Christiano et al., , 2017, Ziegler et al., , 2019, Bai et al., , 2022, Bakker et al., , 2022, Ouyang et al., , 2022], they often suffer from various model misspecifications (see also Table 1 for a summary):
@@
    Preference model misspecification. Most reward-based algorithms rely on the Bradley-Terry [BT, Bradley and Terry, , 1952] preference model (see Equation 1). However, this model entails various unrealistic assumptions on human preference, including transitivity, context-independence and perfect relationality, which are likely violated based on empirical evidence [May, , 1954, Tversky, , 1969, Gardner, , 1970, Agranov and Ortoleva, , 2015, Michaud et al., , 2020, Milano et al., , 2021, Lindner and El-Assady, , 2022]. While some preference-based algorithms impose more general preference model (GPM) assumptions [see e.g., Zhang et al., 2024d, ], their effectiveness still depends on correct model specification.
@@
    Reward model misspecification. Under the BT model assumption, classical reward-based algorithms first estimate the reward function from human preference data and then apply RL algorithms such as the proximal policy optimization [PPO, Schulman et al., , 2017] to derive the optimal policy. However, policy learning through RL is highly sensitive to the estimated reward. Misspecifying the reward can lead to reward hacking [Skalse et al., , 2022, Laidlaw et al., , 2024] and misguide policy learning [Kaufmann et al., , 2023, Zheng et al., , 2023, Chen et al., , 2024].
@@
    Reference policy misspecification. To alleviate misspecification of the reward, recent algorithms based on direct preference optimization [DPO, Rafailov et al., , 2023] propose to express the reward in closed form using the reference policy for policy learning. However, these algorithms are sensitive to the specification of reference policy [Liu et al., 2024b, , Gorbatovski et al., , 2024, Xu et al., , 2024].
@@
  • We propose a robust and efficient estimator for preference evaluation, i.e., evaluating the probability of a target policy being preferred over the reference policy. The proposed preference estimator achieves two desirable properties: (i) double robustness (Corollary 5.6) – it converges to the true preference probability when either the preference model or the reference policy is correctly specified, and (ii) semi-parametric efficiency (Corollary 5.7) – it attains the smallest mean squared error (MSE) among all regular and asymptotically linear estimators [Newey, , 1990, Tsiatis, , 2006].
  • Leveraging this preference estimator, we further develop a preference optimization algorithm for LLM fine-tuning. The proposed algorithm maintains double robustness (Corollary (5.9)) and remains consistent even when the BT model assumption is violated (Theorem 5.8). Meanwhile, when the BT model assumption holds, its suboptimality gap is less sensitive to the reward model and reference policy compared to PPO- or DPO-based algorithms, and is likely smaller than that of both types of algorithms (Theorem 5.10).
@@ 2 Related Works
Reward-based RLHF.
Reward-based algorithms assume the existence of a latent utility or reward function that determines human preferences, estimate the reward function from the data and apply RL for policy learning. Recent research has focused on addressing practical challenges such as reward hacking and model-collapse. These issues arise due to PPO’s sensitivity to reward specification, gradient clipping thresholds, and the tuning parameter controlling Kullback–Leibler (KL)-divergence regularization [Engstrom et al., , 2019, Zheng et al., , 2023, Xiao et al., , 2024]. Existing approaches to these challenges fall into three categories: (i) The first category focuses on improving the reward learning algorithm to obtain more accurate reward functions [Li et al., 2023b, , Chan et al., , 2024, Gao et al., , 2024, Liu et al., 2024a, , Fu et al., , 2025, Xiao et al., , 2025, Ye et al., , 2025]. (ii) The second category develops better policy learning algorithms using the estimated reward function [Wu et al., 2024a, , Zhang et al., 2024a, , Shao et al., , 2024, Hu, , 2025, Liu et al., 2025b, , Yu et al., , 2025, Yuan et al., , 2025]. (iii) The third category is DPO-based, which bypasses reward learning entirely and directly optimizes policies under the BT model assumption [Zhao et al., , 2022, Wang et al., 2023a, , Ethayarajh et al., , 2024, Song et al., , 2024, Tang et al., , 2024].
@@
Preference-based algorithms do not assume the existence of a latent reward function at all; instead, they search the optimal policy that maximizes the alignment with human preferences [see e.g., Azar et al., , 2024]. In particular, there is a growing line of research that adopts the Nash learning from human feedback [NLHF, Munos et al., , 2023] framework, which formulates the alignment problem as a two-player constant-sum game and solves for policies that achieve the Nash equilibrium [Calandriello et al., , 2024, Rosset et al., , 2024, Swamy et al., , 2024, Wu et al., 2024b, , Ye et al., , 2024, Zhang et al., 2024c, , Liu et al., 2025a, ]. Beyond NLHF, Wang et al., 2023b develops a Bayesian approach for alignment, whereas Hong et al., [2024] and Zhang et al., 2024d propose energy-based and general preference models to relax the BT model assumption.
@@
Our proposal belongs to this class of preference-based methods. In particular, the proposed algorithm is most closely related to the identity preference optimization (IPO) algorithm proposed by Azar et al., [2024], as both maximizing the same objective function in the population level (see Section 4 for the objective). However, unlike IPO, our proposed method is robust to misspecifications of the reference policy. Similarly, compared to Zhang et al., 2024d , the proposed algorithm is more robust to the misspecification of the preference model. Finally, our work differs from NLHF in its primary focus: we study robust and statistically efficient preference estimation from data, rather than developing computationally efficient algorithms to solve the Nash equilibrium.
Doubly robust methods.
@@
DR has been extensively studied in statistics, econometrics and machine learning. These methods originate from the missing data and causal inference literature [see e.g., Robins et al., , 1994, Scharfstein et al., , 1999]. To illustrate these methods, consider the fundamental causal inference problem of estimating the average treatment effect (ATE) – the difference in the mean outcome between a newly-developed treatment policy and a baseline policy for a given patient population. DR first estimates two models from the data: (i) a propensity score model (similar to the reference policy in LLMs) that characterizes the treatment assignment mechanism and (ii) an outcome regression model (similar to the reward function) that specifies the conditional mean function of a patient’s outcome. It then employs both models to construct the ATE estimator, whose consistency requires only one of the models to be correct. Furthermore, when both models are correct, the resulting estimator is semiparametrically efficient [Bang and Robins, , 2005]. The favorable statistical properties of these methods have led to extensive follow-up research [see e.g., Tan, , 2010, Tsiatis et al., , 2011, Imai and Ratkovic, , 2014, Vermeulen and Vansteelandt, , 2015, Kennedy et al., , 2017, Robins et al., , 2017, Wager and Athey, , 2018, Wang and Tchetgen Tchetgen, , 2018, Künzel et al., , 2019, Oprescu et al., , 2019, Shi et al., , 2019, Fulcher et al., , 2020, Farrell et al., , 2021, Nie and Wager, , 2021, Li et al., , 2022, Cui et al., , 2023, Han et al., , 2023, Kennedy, , 2023, Li et al., 2023a, , Cui et al., , 2024, Wang and Shah, , 2024, Wang and Han, , 2024]. A seminal extension appears in Chernozhukov et al., [2018], which proposes to learn both the propensity score and outcome regression models using machine learning methods to deal with complex data structures with high-dimensional covariates, texts or images.
@@
Beyond treatment effect estimation in causal inference, doubly robust methods have been widely applied to a broad range of other problems, including the estimation and evaluation of optimal (dynamic) treatment regimes [Robins, , 2004, Zhang et al., , 2012, 2013, Schulte et al., , 2015, Luedtke and Van Der Laan, , 2016, Fan et al., , 2017, Jiang et al., , 2017, Song et al., , 2017, Shi et al., , 2018], conditional independence testing [Zhang et al., , 2019, Shah and Peters, , 2020, Shi et al., , 2021, Quinzan et al., , 2023, Zhang et al., 2024b, ], offline policy learning [Dudík et al., , 2014, Kallus and Uehara, 2020b, , Uehara et al., 2020b, , Liao et al., , 2022, Shi et al., 2024a, ] and off-policy evaluation [OPE, Jiang and Li, , 2016, Thomas and Brunskill, , 2016, Farajtabar et al., , 2018, Kallus and Zhou, , 2018, Bibaut et al., , 2019, Tang et al., , 2020, Kallus and Uehara, 2020a, , Su et al., , 2020, Uehara et al., 2020a, , Shi et al., , 2022, Xie et al., , 2023, Xu et al., , 2023, Cao and Zhou, , 2024, Shi et al., 2024b, ].
@@
RLHF Reward-based PPO-type
DPO-type
Preference-based IPO -
GPM -
@@
3 RLHF Preliminaries
Data generating process.
Assume we are given a dataset $\mathcal{D}$, consisting of $n$ i.i.d. tuples of the form $(X,Y^{(1)},Y^{(2)},Z)$. Each of these tuples is generated as follows: Given a prompt $X$, two independent responses $(Y^{(1)},Y^{(2)})$ are generated under a reference policy $\pi_{\textrm{ref}}$ such that $Y^{(1)},Y^{(2)}\sim\pi_{\textrm{ref}}(\cdot|X)$. These data $(X,Y^{(1)},Y^{(2)})$ are then shown to a human expert, who provides a binary preference $Z=\mathbb{I}(Y^{(1)}\succ Y^{(2)})$ where $Y^{(1)}\succ Y^{(2)}$ indicates that the first response is preferred, and $\mathbb{I}(\bullet)$ denotes the indicator function. Additionally, let $g^{*}$ denote the preference function such that $g^{*}(X,Y^{(1)},Y^{(2)})=\mathbb{P}(Y^{(1)}\succ Y^{(2)}|X)$ determines the probability of $Y^{(1)}$ being favored over $Y^{(2)}$ conditional on $X$.
We remark that the reference policy $\pi_{\textrm{ref}}$ is not always known. For instance, the responses might be generated by an LLM different from the target model that we wish to fine-tune [Bai et al., , 2022]. Furthermore, the responses might be produced by a heterogeneous set of models rather than a single model [Stiennon et al., , 2020].
BT model.
As commented in Section 2, most existing reward-based RLHF algorithms impose the BT model assumption, which requires the preference function $g^{*}$ to take the following form,
@@ 3 RLHF Preliminaries
$J(\pi)=\mathbb{E}[\mathbb{E}_{y\sim\pi(\cdot|X)}r^{*}(y,X)],$ (2)
@@
PPO-based approaches.
PPO-based algorithms proceed in two steps. In the first step, they compute an estimated reward function $\widehat{r}$ using maximum likelihood estimation or empirical risk minimization. In the second step, they learn $\pi^{*}$ by maximizing
$\mathbb{E}_{X\sim\mathcal{D},\,y\sim\pi(\cdot\mid X)}\left[\widehat{r}(y,X)\right]-\beta\,{D}_{\mathrm{KL}}\left[\pi(y\mid X)\,\|\,\pi_{\mathrm{ref}}(y\mid X)\right],$ (3)
over $\pi\in\Pi$ (e.g., a transformer-based policy class), where the expectation is taken over prompts $X$ from the empirical data distribution and responses $y$ from a target policy $\pi$, $D_{\mathrm{KL}}$ denotes the KL divergence measure between the target and reference policies, and the tuning parameter $\beta>0$ controls the degree to which $\pi$ is allowed to deviate from $\pi_{\textrm{ref}}$. The KL regularization term in (3) encourages the learned policy to stay close to $\pi_{\textrm{ref}}$, in order to mitigate over-fitting and prevent the learned policy from collapsing to a narrow set of high-reward responses [Zheng et al., , 2023].
DPO-based approaches.
DPO-based algorithms are motivated by the fact that the argmax to (3) (denoted by $\widehat{\pi}$) can be represented in closed-form using the estimated reward $\widehat{r}$. This in turn yields the following closed-form expression for $\widehat{r}$,
@@ 4 Double Robust Preference Evaluation and Optimization
This section introduces the proposed doubly robust approach; see Figure 1 for a visualization. Different from these reward-based algorithms discussed in Section 3, we adopt a preference-based approach that searches the optimal policy by maximizing its total preference. Specifically, given a target policy $\pi$, its total preference over the reference policy [Azar et al., , 2024] is defined by
$p^{*}(\pi):=\mathbb{P}(\pi\succ\pi_{\textrm{ref}})=\mathbb{E}[\mathbb{E}_{y\sim\pi(\cdot|X),y^{\prime}\sim\pi_{\textrm{ref}}(\cdot|X)}g^{*}(X,y,y^{\prime})],$
@@
$p^{*}(\pi)=\frac{1}{2}\sum_{a=1}^{2}\mathbb{E}[\mathbb{E}_{y\sim\pi(\cdot|X)}g^{*}(X,y,Y^{(a)})].$ (5)
For preference evaluation, our goal is to accurately estimate $p^{*}(\pi)$ for a given target policy $\pi$ from the dataset $\mathcal{D}$. In the following, we first introduce two baseline estimators: a direct method (DM) estimator and an importance sampling (IS) estimator, where the names are borrowed from the OPE literature [see e.g., Uehara et al., , 2022]. We next introduce our proposed DR estimator, which combines both DM and IS for efficient and robust preference evaluation.
DM estimator.
The direct method estimator is motivated by (5). It proceeds by first estimating $g^{*}$ and then plugging the estimated $g^{*}$ (denoted by $\widehat{g}$) into (5) to construct the estimator,
$\widehat{p}_{\textrm{DM}}(\pi)=\frac{1}{2}\mathbb{E}_{X\sim\mathcal{D},y\sim\pi(\cdot|X)}[\widehat{g}(X,y,Y^{(1)})+\widehat{g}(X,y,Y^{(2)})],$ (6)
@@
When an external preference model is available, it can be used directly as $\widehat{g}$, as in [Munos et al., , 2023]. Otherwise, $g^{*}$ can be estimated from the data $\mathcal{D}$. For instance, under the BT model assumption, one can estimate the reward function $r^{*}$ and plug the estimator into (1) to derive $\widehat{g}$. Alternatively, one can employ more general preference models that do not rely on the BT model.
IS estimator.
The second baseline estimator is the IS estimator, which is motivated by the following lemma that expresses $p^{*}(\pi)$ using the IS ratio $w(y,x)=\pi(y|x)/\pi_{\textrm{ref}}(y|x)$.
Lemma 4.1.
Assume $w(y,x)<\infty$ for any $x$, $y$. Then
$p^{*}(\pi)=\frac{1}{2}\mathbb{E}[w(Y^{(1)},X)Z+w(Y^{(2)},X)(1-Z)].$
The proof of Lemma 4.1 is straightforward. It follows directly from the symmetry of pairwise comparisons where the preference can be equivalently expressed using either $g^{*}(X,y,y^{\prime})$ or $1-g^{*}(X,y^{\prime},y)$, and an application of the change-of-measure theorem (see Appendix A.1).
@@
where $\widehat{\pi}_{\textrm{ref}}$ denotes an estimated reference policy. If $\pi_{\textrm{ref}}$ is known, we can directly use the oracle reference policy. Otherwise, for some external datasets [e.g., Bai et al., , 2022], well-trained reference models are available and can be used as $\widehat{\pi}_{\textrm{ref}}$. Finally, when no such external model is available and $\pi_{\textrm{ref}}$ is unknown, we estimate it from the observed data tuples $(X,Y^{(1)},Y^{(2)})$ using supervised fine-tuning (SFT).
DR estimator.
A closer look at Equations (6) and (7) reveals that the DM and IS estimators’ consistencies depend crucially on the correct specification of the preference function and the reference policy. We next introduce our proposed DR estimator, which is more robust to misspecifications in these models. It relies on the following estimating function $\psi(X,Y^{(1)},Y^{(2)},Z;\pi,\widehat{\pi}_{\textrm{ref}},\widehat{g})$, defined as
$\begin{split}\frac{1}{2}\sum_{a=1}^{2}\mathbb{E}_{y\sim\pi(\cdot|X)}[\widehat{g}(X,y,Y^{(a)})]+\frac{1}{2}\sum_{a=1}^{2}(-1)^{a-1}\frac{\pi(Y^{(a)}|X)}{\widehat{\pi}_{\text{ref}}(Y^{(a)}|X)}[Z-\widehat{g}(X,Y^{(1)},Y^{(2)})].\end{split}$ (8)

### v2

@@ 1 Introduction
Recent advances in large language models (LLMs) have revolutionized various natural language processing tasks, ranging from text generation to human-AI conversation and more complex reasoning tasks [1, 2, 3]. These models are typically trained in two stages. In the pre-training stage, LLMs learn general linguistic patterns and commonsense knowledge from vast, unlabeled text data through autoregressive next-token prediction. However, pretrained models face a critical objective mismatch: while they are optimized for token prediction, real-world deployment requires alignment with complex human values such as helpfulness, honesty and harmlessness [4]. This mismatch calls for an additional post-training stage, aiming at better aligning these pre-trained models with human preference.
The paper studies reinforcement learning from human feedback (RLHF), a post-training paradigm that adapts pre-trained models through reinforcement learning [RL, 5]. The RLHF literature has rapidly expanded in recent years, where existing algorithms can be broadly categorized as reward-based or preference-based (Section 2 for a review). While demonstrating remarkable success in domains including robotics control, video games, and LLMs fine-tuning [see e.g., 6, 7, 8, 9, 10], they often suffer from various model misspecifications (see also Table 1 for a summary):
@@
    Preference model misspecification. Most reward-based algorithms rely on the Bradley-Terry [BT, 11] preference model (see Equation 1). However, this model entails various unrealistic assumptions on human preference, including transitivity, context-independence and perfect relationality, which are likely violated based on empirical evidence [12, 13, 14, 15, 16, 17, 18]. While some preference-based algorithms impose more general preference model (GPM) assumptions [see e.g., 19], their effectiveness still depends on correct model specification.
@@
    Reward model misspecification. Under the BT model assumption, classical reward-based algorithms first estimate the reward function from human preference data and then apply RL algorithms such as the proximal policy optimization [PPO, 20] to derive the optimal policy. However, policy learning through RL is highly sensitive to the estimated reward. Misspecifying the reward can lead to reward hacking [21, 22] and misguide policy learning [23, 24, 25].
@@
    Reference policy misspecification. To alleviate misspecification of the reward, recent algorithms based on direct preference optimization [DPO, 26] propose to express the reward in closed form using the reference policy for policy learning. However, these algorithms are sensitive to the specification of reference policy [27, 28, 29].
@@
  • We propose a robust and efficient estimator for preference evaluation, i.e., evaluating the probability of a target policy being preferred over the reference policy. The proposed preference estimator achieves two desirable properties: (i) double robustness (Corollary 3) – it converges to the true preference probability when either the preference model or the reference policy is correctly specified, and (ii) semi-parametric efficiency (Corollary 4) – it attains the smallest mean squared error (MSE) among all regular and asymptotically linear estimators [30, 31].
  • Leveraging this preference estimator, we further develop a preference optimization algorithm for LLM fine-tuning. The proposed algorithm maintains double robustness (Corollary (6)) and remains consistent even when the BT model assumption is violated (Theorem 5). Meanwhile, when the BT model assumption holds, its suboptimality gap is likely smaller than that of PPO- or DPO-based algorithms (Theorem 7).
@@ 2 Related Works
Reward-based RLHF. Reward-based algorithms assume the existence of a latent utility or reward function that determines human preferences, estimate the reward function from the data and apply RL for policy learning. Recent research has focused on addressing practical challenges such as reward hacking and model-collapse. These issues arise due to PPO’s sensitivity to reward specification, gradient clipping thresholds, and the tuning parameter controlling Kullback–Leibler (KL)-divergence regularization [32, 24, 33]. Existing approaches to these challenges fall into three categories: (i) The first category focuses on improving the reward learning algorithm to obtain more accurate reward functions [34, 35, 36, 37, 38, 39, 40]. (ii) The second category develops better policy learning algorithms using the estimated reward function [41, 42, 43, 44, 45, 46, 47]. (iii) The third category is DPO-based, which bypasses reward learning entirely and directly optimizes policies under the BT model assumption [48, 49, 50, 51, 52]. Recent studies have developed robust variants of DPO to handle pairwise noise where preference labels in the training data may be flipped [53, 54, 55].
@@
Preference-based RLHF. Preference-based algorithms do not assume the existence of a latent reward function at all; instead, they search the optimal policy that maximizes the alignment with human preferences [see e.g., 56]. In particular, there is a growing line of research that adopts the Nash learning from human feedback [NLHF, 57] framework, which formulates the alignment problem as a two-player constant-sum game and solves for policies that achieve the Nash equilibrium [58, 59, 60, 61, 62, 63, 64]. Beyond NLHF, [65] develops a Bayesian approach for alignment, whereas [66] and [19] propose energy-based and general preference models to relax the BT model assumption.
@@
Our proposal belongs to this class of preference-based methods. In particular, the proposed algorithm is most closely related to the identity preference optimization (IPO) algorithm proposed by Azar et al. [56], as both maximizing the same objective function in the population level (see Section 4 for the objective). However, unlike IPO, our proposed method is robust to misspecifications of the reference policy. Similarly, compared to [19], the proposed algorithm is more robust to the misspecification of the preference model. Finally, our work differs from NLHF in its primary focus: we study robust and statistically efficient preference estimation from data, rather than developing computationally efficient algorithms to solve the Nash equilibrium.
@@
Doubly robust methods. DR has been extensively studied in statistics, econometrics and machine learning. These methods originate from the missing data and causal inference literature [see e.g., 67, 68]. To illustrate these methods, consider the fundamental causal inference problem of estimating the average treatment effect (ATE) – the difference in the mean outcome between a newly-developed treatment policy and a baseline policy for a given patient population. DR first estimates two models from the data: (i) a propensity score model (similar to the reference policy in LLMs) that characterizes the treatment assignment mechanism and (ii) an outcome regression model (similar to the reward function) that specifies the conditional mean function of a patient’s outcome. It then employs both models to construct the ATE estimator, whose consistency requires only one of the models to be correct. Furthermore, when both models are correct, the resulting estimator is semiparametrically efficient [69]. These methods’ favorable statistical properties have led to extensive follow-up research [see e.g., 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92]. A seminal extension appears in Chernozhukov et al. [93], which proposes to learn both the propensity score and outcome regression models using machine learning methods to deal with complex data structures with high-dimensional covariates, texts or images.
@@
Beyond treatment effect estimation in causal inference, doubly robust methods have been widely applied to a broad range of other problems, including the estimation and evaluation of optimal (dynamic) treatment regimes [94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105], conditional independence testing [106, 107, 108, 109, 110, 111], offline policy learning [112, 113, 114, 115, 116, 117] and off-policy evaluation [OPE, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139].
@@
RLHF Reward-based PPO-based
DPO-based
Preference-based IPO [56] -
GPM [19] -
@@
3 RLHF Preliminaries: Data, Modeling and Baseline Algorithms
Data generating process. Assume we are given a dataset $\mathcal{D}$, consisting of $n$ i.i.d. tuples of the form $(X,Y^{(1)},Y^{(2)},Z)$. Each of these tuples is generated as follows: Given a prompt $X$, two independent responses $(Y^{(1)},Y^{(2)})$ are generated under a reference policy $\pi_{\textrm{ref}}$ such that $Y^{(1)},Y^{(2)}\sim\pi_{\textrm{ref}}(\bullet|X)$. These data $(X,Y^{(1)},Y^{(2)})$ are then shown to a human expert, who provides a binary preference $Z=\mathbb{I}(Y^{(1)}\succ Y^{(2)})$ where $Y^{(1)}\succ Y^{(2)}$ indicates that the first response is preferred, and $\mathbb{I}(\bullet)$ denotes the indicator function. Additionally, let $g^{*}$ denote the preference function such that $g^{*}(X,Y^{(1)},Y^{(2)})=\mathbb{P}(Y^{(1)}\succ Y^{(2)}|X)$ determines the probability of $Y^{(1)}$ being favored over $Y^{(2)}$ conditional on $X$.
We remark that the reference policy $\pi_{\textrm{ref}}$ is not always known. For instance, the responses might be generated by an LLM different from the target model that we wish to fine-tune [8]. Furthermore, the responses might be produced by a heterogeneous set of models rather than a single model [140, 141, 142].
BT model. As commented in Section 2, most existing reward-based RLHF algorithms impose the BT model assumption, which requires the preference function $g^{*}$ to take the following form,
@@ 3 RLHF Preliminaries: Data, Modeling and Baseline Algorithms
$J(\pi)=\mathbb{E}[\mathbb{E}_{y\sim\pi(\bullet|X)}r^{*}(y,X)],$ (2)
@@
PPO-based approaches. PPO-based algorithms proceed in two steps. In the first step, they compute an estimated reward function $\widehat{r}$ using maximum likelihood estimation or empirical risk minimization. In the second step, they learn $\pi^{*}$ by maximizing
$\mathbb{E}_{X\sim\mathcal{D},\,y\sim\pi(\bullet\mid X)}\left[\widehat{r}(y,X)\right]-\beta\,{D}_{\mathrm{KL}}\left[\pi(y\mid X)\,\|\,\pi_{\mathrm{ref}}(y\mid X)\right],$ (3)
over $\pi\in\Pi$ (e.g., a transformer-based policy class), where the expectation is taken over prompts $X$ from the empirical data distribution and responses $y$ from a target policy $\pi$, $D_{\mathrm{KL}}$ denotes the KL divergence measure between the target and reference policies, and the tuning parameter $\beta>0$ controls the degree to which $\pi$ is allowed to deviate from $\pi_{\textrm{ref}}$. The KL regularization term in (3) encourages the learned policy to stay close to $\pi_{\textrm{ref}}$, in order to mitigate over-fitting and prevent the learned policy from collapsing to a narrow set of high-reward responses [24].
DPO-based approaches. DPO-based algorithms are motivated by the fact that the argmax to (3) (denoted by $\widehat{\pi}$) can be represented in closed-form using the estimated reward $\widehat{r}$. This in turn yields the following closed-form expression for $\widehat{r}$,
@@ 4 Double Robust Preference Evaluation and Optimization
This section introduces the proposed doubly robust approach; see Figure 1 for a visualization. Different from these reward-based algorithms discussed in Section 3, we adopt a preference-based approach that searches the optimal policy by maximizing its total preference. Specifically, given a target policy $\pi$, its total preference over the reference policy [56] is defined by
$p^{*}(\pi):=\mathbb{P}(\pi\succ\pi_{\textrm{ref}})=\mathbb{E}[\mathbb{E}_{y\sim\pi(\bullet|X),y^{\prime}\sim\pi_{\textrm{ref}}(\bullet|X)}g^{*}(X,y,y^{\prime})],$
@@
$p^{*}(\pi)=\frac{1}{2}\sum_{a=1}^{2}\mathbb{E}[\mathbb{E}_{y\sim\pi(\bullet|X)}g^{*}(X,y,Y^{(a)})].$ (5)
For preference evaluation, our goal is to accurately estimate $p^{*}(\pi)$ for a given target policy $\pi$ from the dataset $\mathcal{D}$. In the following, we first introduce two baseline estimators: a direct method (DM) estimator and an importance sampling (IS) estimator, where the names are borrowed from the OPE literature [see e.g., 143]. We next introduce our proposed DR estimator, which combines both DM and IS for efficient and robust preference evaluation.
DM estimator. The direct method estimator is motivated by (5). It proceeds by first estimating $g^{*}$ and then plugging the estimated $g^{*}$ (denoted by $\widehat{g}$) into (5) to construct the estimator,
$\widehat{p}_{\textrm{DM}}(\pi)=\frac{1}{2}\mathbb{E}_{X\sim\mathcal{D},y\sim\pi(\bullet|X)}[\widehat{g}(X,y,Y^{(1)})+\widehat{g}(X,y,Y^{(2)})],$ (6)
@@
When an external preference model is available, it can be used directly as $\widehat{g}$, as in [57]. Otherwise, $g^{*}$ can be estimated from the data $\mathcal{D}$. For instance, under the BT model assumption, one can estimate the reward function $r^{*}$ and plug the estimator into (1) to derive $\widehat{g}$. Alternatively, one can employ more general preference models that do not rely on the BT model.
IS estimator. The second baseline estimator is the IS estimator, which is motivated by the following lemma that expresses $p^{*}(\pi)$ using the IS ratio $w(y,x)=\pi(y|x)/\pi_{\textrm{ref}}(y|x)$.
Lemma 1.
Assume $w(y,x)<\infty$ for any $x$, $y$. Then $p^{*}(\pi)=\frac{1}{2}\mathbb{E}[w(Y^{(1)},X)Z+w(Y^{(2)},X)(1-Z)]$.
The proof of Lemma 1 is straightforward. It follows directly from the symmetry of pairwise comparisons where the preference can be equivalently expressed using either $g^{*}(X,y,y^{\prime})$ or $1-g^{*}(X,y^{\prime},y)$, and an application of the change-of-measure theorem (see Appendix A.1).
@@
where $\widehat{\pi}_{\textrm{ref}}$ denotes an estimated reference policy. If $\pi_{\textrm{ref}}$ is known, we can directly use the oracle reference policy. Otherwise, for some external datasets [e.g., 8], well-trained reference models are available and can be used as $\widehat{\pi}_{\textrm{ref}}$. Finally, when no such external model is available and $\pi_{\textrm{ref}}$ is unknown, we estimate it from the observed data tuples $(X,Y^{(1)},Y^{(2)})$ using supervised fine-tuning (SFT).
[…]
$\begin{split}\frac{1}{2}\sum_{a=1}^{2}\mathbb{E}_{y\sim\pi(\bullet|X)}[\widehat{g}(X,y,Y^{(a)})]+\frac{1}{2}\sum_{a=1}^{2}(-1)^{a-1}\frac{\pi(Y^{(a)}|X)}{\widehat{\pi}_{\text{ref}}(Y^{(a)}|X)}[Z-\widehat{g}(X,Y^{(1)},Y^{(2)})].\end{split}$ (8)